OpenStudy (anonymous):

(can someone please help me :( ) Solve each system by graphing. Check your answers. 5. y = -x + 4 y + 3x + 12

OpenStudy (dtan5457):

Graphing? allowed to use a graphing calculator, correct?

OpenStudy (anonymous):

yes

OpenStudy (dtan5457):

Haha. Then it's too easy. Make sure both lines have an isolated y so the first one is done. the 2nd one should be y=-3x+12 plug those 2 equations into https://www.desmos.com/calculator solution: 4,0

OpenStudy (danjs):

The solution (if there is one) to 2 linear equations, is the same as the point where the 2 lines intersect.

OpenStudy (anonymous):

wait so the answer will be 4,0? :o @dtan5457

OpenStudy (dtan5457):

actually wait up. the 2nd one is y + 3x + 12=0?

OpenStudy (dtan5457):

if so then the 2nd line should actually be y=-3x-12 the solution is -8,12 sorry. read it wrong.

OpenStudy (anonymous):

oh wait, it's y= 3x + 12 @dtan5457

OpenStudy (dtan5457):

LOL. then we were all wrong. it'll be -2,6. final answer<<

OpenStudy (dtan5457):

i was wondering how a linear equation would be y+3x+12..

OpenStudy (anonymous):

ohhh! yay! :) can you please please please! help me with a few more? I'm stressing over here @dtan5457

OpenStudy (dtan5457):

go ahead.

OpenStudy (anonymous):

yay! thank you! you are like a life saver

OpenStudy (anonymous):

11. Describe the graph of a system of equations that has no solution

OpenStudy (dtan5457):

they will never intersect ^^

OpenStudy (dtan5457):

anything else?

OpenStudy (anonymous):

yess :3

OpenStudy (dtan5457):

no problem. ask away

OpenStudy (anonymous):

yay okay. let me type it ^_^

OpenStudy (anonymous):

-x + y = -13 3x - y = 19

OpenStudy (dtan5457):

solve it graphically, again?

OpenStudy (anonymous):

yes ^_^

OpenStudy (dtan5457):

anytime you solve a system graphically literally the only work you have to do is isolate y. -x + y = -13 add the x so y=-13+x 3x - y = 19 subtract the 3x -y=19-3x, negative y so negate everything to y=-19+3x plug that in your calculator or https://www.desmos.com/calculator intersects at 3,-10<<<your solution

OpenStudy (dtan5457):

questions? more problems?

OpenStudy (anonymous):

yesss :o if that's okay. am I bother you?

OpenStudy (dtan5457):

not really, as long as your learning it too.

OpenStudy (anonymous):

I am ! but I still need helpp with it and you're doing it :). I tend to forget things something :(

OpenStudy (anonymous):

like I'm really suckish at math

OpenStudy (anonymous):

and im trying to boost my grade up before the semester is over cause it's like literally almost over

OpenStudy (dtan5457):

don't worry i feel you. how many problems you got left? if you only got a few i'll guide you through them so you can learn them properly

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }y = \frac{ 7 }{ 3 }x + \frac{ 5 }{ 3 }\] x - 3y = 5

OpenStudy (anonymous):

hold on. let me see how much left

OpenStudy (anonymous):

4 more left

OpenStudy (dtan5457):

alright look it's the same thing i said before, isolate y. are you able to do that for these equations

OpenStudy (anonymous):

I think but can you tell me how again? like I get it but I don't get it :/

OpenStudy (anonymous):

(sorry if I'm annoying D: )

OpenStudy (dtan5457):

you have to make y literally just y=........etc etc 1/3y=7/3x+5/3

OpenStudy (dtan5457):

you have 1/3 y how can you make it just y

OpenStudy (dtan5457):

for solving any system you must get that concept

OpenStudy (dtan5457):

multiply everything by 3 lol

OpenStudy (anonymous):

hold on, I'm looking back and forth

OpenStudy (anonymous):

ohhh

OpenStudy (dtan5457):

tell me what you get

OpenStudy (dtan5457):

are you having trouble with this

OpenStudy (anonymous):

maybe ... ._.

OpenStudy (anonymous):

omg IM SO STUPID :( IM SORRY

OpenStudy (dtan5457):

LOL it's ok i failed plenty of math tests this year

OpenStudy (anonymous):

yeah but I really need to get this done or I'm going to fail my math class :(

OpenStudy (anonymous):

like my whole math class

OpenStudy (anonymous):

Imma cry

OpenStudy (anonymous):

*starts to hyperventilate*

OpenStudy (dtan5457):

if you multiply it by 3 you get y=7x+5

OpenStudy (anonymous):

but what am I multiplyingg D: I forgottt!

OpenStudy (dtan5457):

well look if you have 1/3y you wanna get 1y basically you gonna multiply by 3 cause 1/3 times 3/1=3

OpenStudy (dtan5457):

i think your thinking too hard

OpenStudy (anonymous):

I think I am! all I can think about is getting these assessments done so I can stop stressing and breathe!

OpenStudy (anonymous):

wait so what am I doing with y= 7x + 5? putting it in Desmos?

OpenStudy (anonymous):

*faints*

OpenStudy (dtan5457):

yep. the 2nd one is x - 3y = 5 isolating y is -3y=5-x divide by -3 y=-5/3+x/3

OpenStudy (dtan5457):

put that in the calculator and tell me where it intersects

OpenStudy (anonymous):

I got 0, 5??? o.o

OpenStudy (dtan5457):

you typed it in wrong

OpenStudy (dtan5457):

if anything you have to type it in correctly

OpenStudy (anonymous):

ughhh sfgvkjcnfdjvfkmredsjc,fdcmk

OpenStudy (anonymous):

how do I type it in right?

OpenStudy (dtan5457):

y=-5/3+x/3

OpenStudy (dtan5457):

y=7x+5

OpenStudy (dtan5457):

you know what i mean

OpenStudy (anonymous):

5,0 ? here let me show you how I typed it

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

yea that looks right

OpenStudy (dtan5457):

now do the other one

OpenStudy (anonymous):

so the answer is 5, 0?

OpenStudy (dtan5457):

according to my calculator its -1,-2

OpenStudy (dtan5457):

how did u get 5,0

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

I don't knoww! D:

OpenStudy (anonymous):

how do you see where the answer is?

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

yeah I'm lost. where did you find the answer <__>

OpenStudy (dtan5457):

where it look like it intersect

OpenStudy (dtan5457):

wasup

OpenStudy (anonymous):

<_> ._. >_> :( oh

OpenStudy (anonymous):

I'll write this all down in my notepads so I can remember as we go on

OpenStudy (dtan5457):

awesome next question

OpenStudy (dtan5457):

if they let you use a graphing calculator there should be no reason to get this wrong

OpenStudy (anonymous):

they do but like I don't have oneeee D: and I do online school :(

OpenStudy (dtan5457):

online school? that exists?

OpenStudy (anonymous):

okay let me get the next question

OpenStudy (anonymous):

17. 3x + y = -13 -2x + 5y = 054

OpenStudy (anonymous):

and yes they exist

OpenStudy (anonymous):

and I have a math assessment that's dued tonight >___> which is the one that we are doing <_< that you are helping me with >_> which I'm happy for cause otherwise I'll be crying

OpenStudy (dtan5457):

-2x + 5y = 054 typo or nah

OpenStudy (dtan5457):

is it 54?

OpenStudy (anonymous):

oops